Network Interruption 3_2_10
Normal Operating Conditions

March 2, 2010

This morning a slow leak was discovered in the air conditioning unit in the Network Operations Center (NOC).  At this time, the temperature in the NOC is being maintained at a safe level and we are working with maintenance to keep the AC unit running in its current condition as long as possible.

At 3 p.m., the maintenance staff will shut down the AC unit to begin repairs. We anticipate the temperatures in the NOC to rise to such levels that we will need to shut down many, if not all, of the servers. 

During the repairs the following services will not be available:
Web access to email
Student login servers
S and T drives

We will attempt to keep the faculty and staff servers (email, H drive, logins, J drive) running until 5 p.m. We will try to keep the web servers and redwood running as long as possible.

It is anticipated that the repairs will continue through the night and that we will be able to resume normal operations by morning.

We appreciate your patience.

Posted in: